REU: SCHEDULE FOR WEEKS THREE AND FOUR

Date  YSP  Babai  Sally  May
Monday, July 1  9:30 – 2:30  2:45 – 4:15
Tuesday, July 2  9:30 – 2:30 5:00 – 6:30*  2:45 – 4:15
Wednesday, July 3  9:30 – 2:30  2:45 – 4:15
Thursday, July 4  ——–  ——–  ——–  ——–
Friday, July 5  ——–  1:30 – 4:30  9:30 – 12:00
Monday, July 8  9:30 – 2:30 5:00 – 6:00 2:45 – 4:15
Tuesday, July 9 9:30 – 2:30 5:00 – 6:00* 2:45 – 4:15
Wednesday, July 10 9:30 – 2:30 5:00 – 6:00 2:45 – 4:15
Thursday, July 11 9:30 – 2:30 5:00 – 6:00* 2:45 – 4:15
Friday, July 12  9:30 – 2:00** 5:00 – 6:00 2:45 – 4:15

*Problem sessions (Babai will be away July 1–3) **Staff meeting 12:00 – 2:00 Graduate student sessions will be organized around this schedule.

Here is the schedule in summary form:

Third week:

YSP: 9:30 – 2:30 M – W Babai 5:00 – 6:30 Tues (problem session); 1:30 – 4:30 Friday May 2:45 – 4:15 Tues; 9:30 – 12:00 Friday Sally 2:45 – 4:45 MW

Fourth week:

YSP 9:30 – 2:30 M – F (to 2:00 on Friday) Babai 5:00 – 6:00 M – F (problem sessions TTh) May 2:45 – 4:15 TTh Sally 2:45 – 4:15 MWF